This equation is not in the textbook.
Could some one work this so I can check my son's answer. He got X=-11
would you show the steps for me so i can check how he worked it?
7(x-1)+5=2(3-4x)+7

Found 2 solutions by Lacey020991, checkley71:
Answer by Lacey020991(49)


--Your first step would be to distribute everything.
--Your second step would be to combine like terms
--Your third step out be to add 8x to each side. The one's on the right would cancel out.
--Your fourth step would be to add two to each side. The 2's on the left will cancel out.
--Your fifth step would be to divide by 15.
7(x-1) + 5 = 2(3-4x) + 7
7x - 7 + 5 = 6 - 8x + 7
7x - 2 = 13 - 8x
15x - 2 = 13
15x = 15
x = 1
.
Always check your work:
7(x-1) + 5 = 2(3-4x) + 7
7(1-1) + 5 = 2(3-4*1) + 7
0 + 5 = 6 - 8 + 7
5 = 5
.
I got x = 1. and it checked. I'm not sure how your son got -11. If you need further explaining, just let me know.

Answer by checkley71(8403)

7(X-1)+5=2(3-4X)+7
7X-7+5=6-8X+7
7X+8X=6+7+7-5
15X=15
X=15/15
X=1
PROOF
7(1-1)+5=2(3-4*1)+7
7*0+5=2(3-4)+7
5=2(-1)+7
5=-2+7
5=5